OpenStudy (anonymous):

Simplify x to the fourth power minus 3 times x cubed all over x squared minus 9

OpenStudy (anonymous):

the expression to be simplified is: (x^4-3x^3)/(x^2-9) taking x^3 common out of the numerator and applying the identity a^2-b^2=(a+b)(a-b) in the denominator, we get (x^3)(x-3)/(x+3)(x-3) = (x^3)/(x+3)

OpenStudy (anonymous):

THANKKS